OpenStudy (anonymous):

Evaluate √ x+ 7 when x= 9 A. 4 B. 8 C. 10 D. 16 Help? Please? (: will give a medal

OpenStudy (alexandervonhumboldt2):

ok so sqrt9=3 i hope you know what is 3+7=?

OpenStudy (alexandervonhumboldt2):

if sqrt(x+7)=sqrt(9+7)=sqrt(16)=4 if sqrtx + 7 =then 10